# Federated Behavioural Planes: Explaining the Evolution of Client Behaviour in Federated Learning

### v1

@@ Abstract
Federated Learning (FL), a privacy-aware approach in distributed deep learning environments, enables many clients to collaboratively train a model without sharing sensitive data, thereby reducing privacy risks. However, enabling human trust and control over FL systems requires understanding the evolving behaviour of clients, whether beneficial or detrimental for the training, which still represents a key challenge in the current literature. To address this challenge, we introduce Federated Behavioural Planes (FBPs), a novel method to analyse, visualise, and explain the dynamics of FL systems, showing how clients behave under two different lenses: predictive performance (error behavioural space) and decision-making processes (counterfactual behavioural space). Our experiments demonstrate that FBPs provide informative trajectories describing the evolving states of clients and their contributions to the global model, thereby enabling the identification of clusters of clients with similar behaviours. Leveraging the patterns identified by FBPs, we propose a robust aggregation technique named Federated Behavioural Shields to detect malicious or noisy client models, thereby enhancing security and surpassing the efficacy of existing state-of-the-art FL defense mechanisms.
@@ 1 Introduction
Figure 1: Federated Behavioural Spaces framework makes it possible to visual client behaviour in FL from two perspectives: predictive performance (Error Behavioural Plane) and decision-making processes (Counterfactuals Behavioural Plane). It highlights client trajectories and similarities, offering insights into client interactions and supporting the introduction of a new and effective robust aggregation mechanism with performance that surpasses state-of-the-art baselines.
@@ 2 Background
Federated learning FL systems [20, 3] involve a network of $K\in\mathcal{N}$ clients, coordinated by a central server, which collaboratively train a DL model. Each client $k$ possesses a local and private dataset characterised by a set of $z\in\mathcal{N}$ features $x^{(k)}\in X^{(k)}\subseteq\mathcal{R}^{z}$ and a set of $u\in\mathcal{N}$ class labels $y^{(k)}\in Y^{(k)}\subseteq\{0,1\}^{u}$. In each training round $t\in\mathcal{N}$, each client trains a model $f:X^{(k)}\to Y^{(k)}$ on local data to maximise the likelihood $\mathcal{L}(\theta^{(k)}\mid x^{(k)},y^{(k)})$. Once trained, clients send their local model’s parameters $\theta^{(k)}(t)\in\mathcal{R}^{q},q\in\mathcal{N}$ to a central server which aggregates these parameters using a permutation-invariant aggregation $\oplus:\mathcal{R}^{q\times K}\to\mathcal{R}^{q}$ (such as the mean or median). The server then sends the aggregated model parameters $\theta(t+1)$ back to the clients to start a new training round:
@@
Counterfactual explanations Counterfactual explanations [19] describe a model’s decision-making process by identifying minimal and plausible changes to an observed input’s features that lead to a desired model prediction. In explainable AI, finding counterfactual explanations is framed as an optimisation problem where the objective is to identify, for each sample $x$, the nearest data point $x^{\prime}$ such that the classifier $f(\theta,x^{\prime})$ assigns a desired class label $y^{\prime}$:
@@
Semantic and behavioural spaces Semantic spaces [16, 17] represent the semantics of a model by considering the error vector $e=[\epsilon(f(\theta,x_{i}),y_{i})]_{i=1,\dots,n}$ that a model produces on a set of $n\in\mathcal{N}$ samples, where $\epsilon$ could be the Mean Squared Error, for instance. Given a set of $K$ models, the semantic space contains $n$-dimensional data points $[e_{1},\dots,e_{K}]\in\mathcal{R}^{K\times n}$. Behavioural spaces [18] summarise semantic spaces into lower-dimensional spaces applying a transformation $\psi_{n\to m}:\mathcal{R}^{n}\to\mathcal{R}^{m}$ with $m\ll n$ (typically $m=2$ for most applications) i.e., $\psi_{n\to m}([e_{1},\dots,e_{K}])$.
@@ 3.1 Dynamic behaviour of federated learning
FL is a dynamic process transitioning from a state where all participating clients behave randomly to a state where the behaviour of participating entities becomes coherent (the parameters of client models tend to converge to values which lead to high model accuracy). Thus, FL systems can be effectively analysed through the lens of dynamical systems using tools traditionally employed for such studies, such as differential equations. Defining $b(\theta^{(k)})$ as the evolving state of the behaviour of the client $k$ at time $t$, we introduce two primary forces that influence $b(\theta^{(k)})$: $g(\theta^{(k)},x^{(k)},y^{(k)})$, the local training dynamics, which drives the client towards its local optimum by leveraging information from the local dataset $(x^{(k)},y^{(k)})$; and $b\left(\bigoplus_{k\in K}\theta^{(k)}\right)-b(\theta^{(k)})$, a correction term that periodically aligns $b(\theta^{(k)})$ with the aggregated state of all clients within the federated system. These dynamics can be encapsulated in the following differential equation [^1] For simplicity, we omit the dependency on the time variable $t$ in our notation for variables $b$ and $\theta$.:
@@ 3.4 Counterfactual Behavioural Plane (CBP)
The analysis of a model’s decision-making process is the main research objective of explainable AI. Counterfactual explanations represent one of the most effective techniques as they give insights concerning the position and orientation of decision boundaries. Similar counterfactual explanations indicate models having similar decision boundaries, i.e. models taking decisions using a similar decision-making process. To this end, most differentiable counterfactual generators are trained to model the training data distribution [34, 35], thus potentially providing insights on non identical distributions in clients’ data. To produce these explanations, clients’ predictive models should be concurrently trained with a counterfactual generator. Additional information on the optimisation objective are provided in Appendix A. To obtain the Counterfactual Behavioural Plane (CBP), we first compute the distances between the counterfactual distribution generated by the server using a model of the client $k$ and the distribution of other clients. Then, we apply dimensionality reduction to obtain the CBP.
@@
CBP produces complementary information to EBP as clients which are similar in the CBP might be far away in the EBP (as discussed in Appendix B). Furthermore, as the purpose of CBP is to track clients’ behaviour rather than explaining the model decision to a user, counterfactuals can be generated based on predictive models’ embeddings, instead of input features, concealing sensitive information.
@@ 4 Experiments
This section describes essential information about the experiments. We provide further details in Appendix A.2 such as counterfactual generators, model optimisation, and training setup.
@@ 4.1 Data & task setup
In our experiments, we utilise four datasets: a Synthetic dataset (tabular) we designed to have full control on clients’ data distributions, and thus test our assumptions; the Breast Cancer Wisconsin [36] (tabular); the Diabetes Health Indicator [37] (tabular); and small-MNIST [38] (image), reducing its size by 76% to increase task difficulty and highlight client differences in performance. For all the experiments with the small-MNIST dataset, our approach involves generating counterfactuals at a non-interpretable internal representation level of the model instead of the input space. To reflect the most realistic cross-silo scenario [39], we distribute the training data among various clients such as data are not independent and identically distributed (IID) [40]. This represents the most challenging scenario to detect malicious clients due to the significant variations even among benign clients. Further details on the datasets and non-IID implementation are provided in Appendix A.1. Additional experiments analysing setup characteristics such as window length, local epochs, server validation set size, and differences between non-IID and IID scenarios can be found in Appendix B.
@@ 4.2 Evaluation
Metrics To determine the efficacy of counterfactuals generated through end-to-end training in FL, we measure several key metrics: task accuracy ($\uparrow$ – higher is better); counterfactual validity ($\uparrow$) [19], which checks if the counterfactuals’ labels align with user-provided labels; proximity ($\downarrow$) [35], assessing whether counterfactuals are realistic by their closeness to the training data; and sparsity ($\downarrow$) [19], which quantifies the changes made to the input to generate the counterfactuals. The latter is quantified using Euclidean distance, as counting the number of changes provides less insight on the generated counterfactuals. To evaluate the effectiveness of client-specific adaptation, we analyse the relative change in client proximity between global and client-specific models, expressed as $(P_{global}-P_{local})/P_{global}$. Finally, we measure the task accuracy ($\uparrow$) of the FL system under different attacks and defenses. All metrics are reported as the mean and standard error across five experimental runs with distinct parameter initialization.
Baselines In our experiments, we compare Federated Behavioural Shields with the following state-of-the-art robust aggregation methods: Median [11], Trimmed-mean [11] and Krum [12]. Further information is provided in the Appendix A.7.
Federated Attacks We focus on attacks with realistic assumptions and where additional information outside the typical FL scenario are not available [10, 14, 23]. We test the following attacks: Label-flipping (Data Poisoning) [26, 41], changes each sample’s label to $1-y$ (performed in the context of binary classification); Inverted-loss (Data Poisoning) [25], creates an update that maximises the loss on the local dataset; Crafted-noise (Model Poisoning) [42], adds noise $\mathcal{N}(0,\beta\cdot\sigma(w^{t}))$ to the previous global model $w^{t}$, where $\sigma(w^{t})$ is the standard deviation of $w^{t}$ and $\beta$ is a scale factor set to 1.2; and Inverted-gradient (Model Poisoning) [24, 43], inverts the gradient derived from the server’s previous update, misaligning it with the true gradient. Further details are provided in Appendix A.7.
@@ 5.1 Counterfactuals in FL
Integrating counterfactual generators in FL optimisation does not compromise predictive performance. (Table 1) We compare the models’ accuracy across three distinct settings and three different datasets under non-IID conditions. The results indicate that our model, which involves the concurrent training of both the predictor and the counterfactual generator, achieves performance comparable to that of the predictor alone in FL. In the context of our model, both Local Centralised Learning (CL), a CL performed on individual client data, and FL, implemented across all clients, adhere to privacy standards [2]. However, only FL reaches the performance levels of the CL scenario, which assumes local access to all client data. This result indicates that incorporating counterfactuals during training does not compromise the predictor’s performance, thus supporting their beneficial application without adverse effects on performance.
Counterfactuals generated in FL have similar quality to the ones in CL. (Table 1) Unlike Local CL, FL leverages information from all clients, thereby producing counterfactuals that more accurately reflect non-IID clients’ data. As shown in Table 1, FL achieves higher validity compared to the Local CL approach, indicating that the FL’s counterfactuals better match ground-truth labels. Additionally, FL exhibits lower sparsity, which measures the number of modifications needed to achieve the counterfactuals. Table 1 shows these results using a counterfactual generator we adapted for this scenario (see Appendix A.2), however, similar conclusions are also obtained using out-of-the-box generators [34] (see Appendix B).
@@
Counterfactual generators can be adapted to specific clients. (Figure 2) In our study, we explored the impact of client adaptation on counterfactual generation within FL environments, as detailed in Section 4.2. Client-personalisation is key whenever we need to extract client-specific information. This adaptation can be achieved by training (or fine-tuning) a counterfactual generator on local client’s data (which naturally happens at each round). The effectiveness of client-specific adaptation can be measured by the relative change in client proximity between global and client-specific models, as shown in Figure 2. The figure shows a marked reduction in relative proximity across the three datasets under non-IID conditions (up to 70% in the Synthetic dataset). This suggests that client-specific counterfactuals after adaptation are more representative of individual client datasets, providing unique descriptors of client-specific behaviours in the CBP.
@@ 5.2 Explaining FL training
Figure 3: Client trajectories on Counterfactuals and Error Behavioural Planes for Synthetic, Breast Cancer, and small-MNIST datasets, corresponding to Inverted-loss, Crafted-noise, and Inverted-gradient attacks, respectively. The figure highlights the deviation of the malicious client (red) from honest clients, who tend to cluster together over time.
@@
The trajectories in FBPs enable the identification of different client behaviours during training. (Figure 3) Figure 3 illustrates FBPs’ client trajectories over the last 15 rounds, where we introduced a different attack on each dataset (from left to right: Inverted-loss, Crafted-noise, and Inverted-gradient attacks). Each dataset was configured with five non-IID clients and one attacker (red). In the Synthetic dataset, to highlight the visualization of honest client clusters, the two largest clients were subdivided into three (Client 2,3,4) and two smaller clients (Client 6,7), respectively, forming distinct clusters. Specifically, on the EBP for the Synthetic dataset, the attacker (Client 9) significantly deviates from the trajectories of other clients, aiming to disrupt the server model (S). On the CBP, clusters with similar data distributions converge, indicating similarity in decision boundaries and client training data. In Breast Cancer and small-MNIST, FBPs are also able to give insights on the type of attack. On the Breast Cancer EBP, Client 6 exhibits a random trajectory around the server trajectory, which indicates a Crafted-noise attack. In small-MNIST, both the CBP and EBP reveal that at each round the trajectory of the attacker consistently moves in the opposite direction to that of the server, which indicates an Inverted-gradient attack. These insights might be useful for FL users as they explain the nature of clients’ behaviours, allowing the identification of as various types of attacks. Additional visualizations are available in the Appendix B.
@@
FBPs allow the identification of clusters of clients. (see Figure 3) In the Synthetic dataset, CBPs’s clusters reflect client-specific data distributions (Client 1,2-3-4,5,6-7,8 and Attacker 9). Clients sharing similar data distributions tend to cluster closely in the CBP, which in the Synthetic dataset is structured with adjacent slices in the feature space (for example, Client 1 is positioned between Client 8 and Client 2). Similarly, in Breast Cancer and small-MNIST, two primary clusters are identifiable in the CBP: honest clients and attackers. This clustering capability is crucial for users aiming to comprehend client characteristics without direct access to data or models’ parameters, thereby informing strategic decisions in training a federated model. For instance, one might consider reducing the number of clients with identical data distributions to avoid redundancy and enhance training efficiency. Conversely, identifying distinct clusters in FBPs can be useful to maximise model performance (at the expense of the generalisation ability) by forming a Clustered FL system where independent federated models are trained using a subset of similar clients.

### v2

@@ Abstract
Federated Learning (FL), a privacy-aware approach in distributed deep learning environments, enables many clients to collaboratively train a model without sharing sensitive data, thereby reducing privacy risks. However, enabling human trust and control over FL systems requires understanding the evolving behaviour of clients, whether beneficial or detrimental for the training, which still represents a key challenge in the current literature. To address this challenge, we introduce Federated Behavioural Planes (FBPs), a novel method to analyse, visualise, and explain the dynamics of FL systems, showing how clients behave under two different lenses: predictive performance (error behavioural space) and decision-making processes (counterfactual behavioural space). Our experiments demonstrate that FBPs provide informative trajectories describing the evolving states of clients and their contributions to the global model, thereby enabling the identification of clusters of clients with similar behaviours. Leveraging the patterns identified by FBPs, we propose a robust aggregation technique named Federated Behavioural Shields to detect malicious or noisy client models, thereby enhancing security and surpassing the efficacy of existing state-of-the-art FL defense mechanisms. Our code is publicly available on GitHub [^1] https://github.com/dariofenoglio98/CF_FL.
footnotetext: Author contributions are detailed in the Acknowledgement section.
@@ 1 Introduction
Figure 1: The Federated Behavioural Planes framework enables the visualization of client behaviour in FL from two perspectives: predictive performance (Error Behavioural Plane) and decision-making processes (Counterfactuals Behavioural Plane). It highlights client trajectories and similarities, offering insights into client interactions and supporting the introduction of a new and effective robust aggregation mechanism with performance that surpasses state-of-the-art baselines.
@@ 2 Background
Federated learning.
FL systems [20, 3] involve a network of $K\in\mathcal{N}$ clients, coordinated by a central server, which collaboratively train a DL model. Each client $k$ possesses a local and private dataset characterised by a set of $z\in\mathcal{N}$ features $x^{(k)}\in X^{(k)}\subseteq\mathcal{R}^{z}$ and a set of $u\in\mathcal{N}$ class labels $y^{(k)}\in Y^{(k)}\subseteq\{0,1\}^{u}$. In each training round $t\in\mathcal{N}$, each client trains a model $f:X^{(k)}\to Y^{(k)}$ on local data to maximise the likelihood $\mathcal{L}(\theta^{(k)}\mid x^{(k)},y^{(k)})$. Once trained, clients send their local model’s parameters $\theta^{(k)}(t)\in\mathcal{R}^{q},q\in\mathcal{N}$ to a central server which aggregates these parameters using a permutation-invariant aggregation $\oplus:\mathcal{R}^{q\times K}\to\mathcal{R}^{q}$ (such as the mean or median). The server then sends the aggregated model parameters $\theta(t+1)$ back to the clients to start a new training round:
@@
Counterfactual explanations.
Counterfactual explanations [19] describe a model’s decision-making process by identifying minimal and plausible changes to an observed input’s features that lead to a desired model prediction. In explainable AI, finding counterfactual explanations is framed as an optimisation problem where the objective is to identify, for each sample $x$, the nearest data point $x^{\prime}$ such that the classifier $f(\theta,x^{\prime})$ assigns a desired class label $y^{\prime}$:
@@
Semantic and behavioural spaces.
Semantic spaces [16, 17] represent the semantics of a model by considering the error vector $e=[\epsilon(f(\theta,x_{i}),y_{i})]_{i=1,\dots,n}$ that a model produces on a set of $n\in\mathcal{N}$ samples, where $\epsilon$ could be the Mean Squared Error, for instance. Given a set of $K$ models, the semantic space contains $n$-dimensional data points $[e_{1},\dots,e_{K}]\in\mathcal{R}^{K\times n}$. Behavioural spaces [18] summarise semantic spaces into lower-dimensional spaces applying a transformation $\psi_{n\to m}:\mathcal{R}^{n}\to\mathcal{R}^{m}$ with $m\ll n$ (typically $m=2$ for most applications) i.e., $\psi_{n\to m}([e_{1},\dots,e_{K}])$.
@@ 3.1 Dynamic behaviour of federated learning
FL is a dynamic process transitioning from a state where all participating clients behave randomly to a state where the behaviour of participating entities becomes coherent (the parameters of client models tend to converge to values which lead to high model accuracy). Thus, FL systems can be effectively analysed through the lens of dynamical systems using tools traditionally employed for such studies, such as differential equations. Defining $b(\theta^{(k)})$ as the evolving state of the behaviour of the client $k$ at time $t$, we introduce two primary forces that influence $b(\theta^{(k)})$: $g(\theta^{(k)},x^{(k)},y^{(k)})$, the local training dynamics, which drives the client towards its local optimum by leveraging information from the local dataset $(x^{(k)},y^{(k)})$; and $b\left(\bigoplus_{k\in K}\theta^{(k)}\right)-b(\theta^{(k)})$, a correction term that periodically aligns $b(\theta^{(k)})$ with the aggregated state of all clients within the federated system. These dynamics can be encapsulated in the following differential equation [^2] For simplicity, we omit the dependency on the time variable $t$ in our notation for variables $b$ and $\theta$., which describes how client behaviours evolve during training and are influenced by internal forces within the FL system:
@@ 3.4 Counterfactual Behavioural Plane (CBP)
The analysis of a model’s decision-making process is the main research objective of explainable AI. Counterfactual explanations represent one of the most effective techniques as they give insights concerning the position and orientation of decision boundaries. Similar counterfactual explanations indicate models having similar decision boundaries, i.e. models taking decisions using a similar decision-making process. To this end, most differentiable counterfactual generators are trained to model the training data distribution [34, 35], thus potentially providing insights on non identical distributions in clients’ data. To produce these explanations, clients’ predictive models should be concurrently trained with a counterfactual generator. Additional information on the optimisation objective are provided in Appendix A.2, A.3. To obtain the Counterfactual Behavioural Plane (CBP), we first compute the distances between the counterfactual distribution generated by the server using a model of the client $k$ and the distribution of other clients. Then, we apply dimensionality reduction to obtain the CBP.
@@
CBP produces complementary information to EBP as clients which are similar in the CBP might be far away in the EBP (as discussed in Appendix B.4). Furthermore, as the purpose of CBP is to track clients’ behaviour rather than explaining the model decision to a user, counterfactuals can be generated based on predictive models’ embeddings, instead of input features, concealing sensitive information.
@@ 4 Experiments
This section describes essential information about the experiments. Further details on model configuration, training setup, and computational cost are presented in Appendices A.2, A.4, and A.6, respectively.
@@ 4.1 Data & task setup
In our experiments, we utilise four datasets: a Synthetic dataset (tabular) we designed to have full control on clients’ data distributions, and thus test our assumptions; the Breast Cancer Wisconsin [36] (tabular); the Diabetes Health Indicator [37] (tabular); small-MNIST [38] (image); and small-CIFAR-10 [39] (image), reducing its size by 76% to increase task difficulty and highlight client differences in performance. For all the experiments with the small-MNIST dataset, our approach involves generating counterfactuals at a non-interpretable internal representation level of the model instead of the input space. To reflect the most realistic cross-silo scenario [40], we distribute the training data among various clients such as data are not independent and identically distributed (IID) [41]. This represents the most challenging scenario to detect malicious clients due to the significant variations even among benign clients. Further details on the datasets and non-IID implementation are provided in Appendix A.1. Additional experiments analysing setup characteristics such as window length, local epochs, server validation set size, and differences between non-IID and IID scenarios can be found in Appendix B.
@@ 4.2 Evaluation
Metrics.
To determine the efficacy of counterfactuals generated through end-to-end training in FL, we measure several key metrics: task accuracy ($\uparrow$ – higher is better); counterfactual validity ($\uparrow$) [19], which checks if the counterfactuals’ labels align with user-provided labels; proximity ($\downarrow$) [35], assessing the realism of counterfactuals by their closeness to the training data (distance between the counterfactual and the closest data point in the training set with the same label); and sparsity ($\downarrow$) [19], which quantifies the changes made to the input to generate the counterfactuals (number of features changed between the initial sample and the counterfactual). The latter is quantified using Euclidean distance, as counting the number of changes provides less insight on the generated counterfactuals. To evaluate the effectiveness of client-specific adaptation, we analyse the relative change in client proximity between global and client-specific models, expressed as $(P_{global}-P_{local})/P_{global}$. Finally, we measure the task accuracy ($\uparrow$) of the FL system under different attacks and defenses. All metrics are reported as the mean and standard error across five experimental runs with distinct parameter initialization.
Baselines.
In our experiments, we compare Federated Behavioural Shields with the following state-of-the-art robust aggregation methods: Median [11], Trimmed-mean [11], Krum [12], and RFA [42]. Further information is provided in the Appendix A.7.
Federated Attacks.
We focus on attacks with realistic assumptions and where additional information outside the typical FL scenario are not available [10, 14, 23]. We test the following attacks: Label-flipping (Data Poisoning) [26, 43], changes each sample’s label to $1-y$ (performed in the context of binary classification); Inverted-loss (Data Poisoning) [25], creates an update that maximises the loss on the local dataset; Crafted-noise (Model Poisoning) [44], adds noise $\mathcal{N}(0,\beta\cdot\sigma(w^{t}))$ to the previous global model $w^{t}$, where $\sigma(w^{t})$ is the standard deviation of $w^{t}$ and $\beta$ is a scale factor set to 1.2; and Inverted-gradient (Model Poisoning) [24, 45], inverts the gradient derived from the server’s previous update, misaligning it with the true gradient. Further details are provided in Appendix A.8.
@@ 5.1 Counterfactuals in FL
Integrating counterfactual generators in FL optimisation does not compromise predictive performance (Table 1). We compared model accuracy across four settings: two centralised learning (CL) and two FL, using three different datasets under non-IID conditions. For FL experiments, we used the traditional FedAvg approach with two variations: predictor-only and predictor with CF generator. The results indicate that our model, which involves the concurrent training of both the predictor and the counterfactual generator, achieves performance comparable to that of the predictor alone in FL. In the context of our model, both Local CL—where each client trains a model on its local data—and FL, implemented across all clients, comply with privacy standards [2]. For Local CL, we report the average accuracy of models trained independently by each client and evaluated on a common test set. However, only FL reaches the performance levels of the CL scenario, which assumes local access to all client data. This result indicates that incorporating counterfactuals during training does not compromise the predictor’s performance, thus supporting their beneficial application without adverse effects on performance.
Counterfactuals generated in FL have similar quality to those in CL (Table 1). Unlike Local CL, FL leverages information from all clients, thereby producing counterfactuals that more accurately reflect non-IID clients’ data. As shown in Table 1, FL achieves higher validity compared to the Local CL approach, indicating that the FL’s counterfactuals better match ground-truth labels. Additionally, FL exhibits lower sparsity, which measures the number of modifications needed to achieve the counterfactuals. Table 1 shows these results using a counterfactual generator we adapted for this scenario (see Appendix A.3), however, similar conclusions are also obtained using out-of-the-box generators [34] (see Appendix B.2).
@@
Counterfactual generators can be adapted to specific clients (Figure 2). In our study, we explored the impact of client adaptation on counterfactual generation within FL environments, as detailed in Section 4.2. Client-personalisation is key whenever we need to extract client-specific information. This adaptation can be achieved by training (or fine-tuning) a counterfactual generator on local client’s data (which naturally happens at each round). The effectiveness of client-specific adaptation can be measured by the relative change in client proximity between global and client-specific models, as shown in Figure 2. The figure shows a marked reduction in relative proximity across the three datasets under non-IID conditions (up to 70% in the Synthetic dataset). This suggests that client-specific counterfactuals after adaptation are more representative of individual client datasets, providing unique descriptors of client-specific behaviours in the CBP.
@@ 5.2 Explaining FL training
Figure 3: Client trajectories on Counterfactuals and Error Behavioural Planes for Synthetic, Breast Cancer, and small-MNIST datasets, corresponding to Inverted-loss, Crafted-noise, and Inverted-gradient attacks, respectively. The figure highlights the deviation of the malicious client (red) from honest clients, who tend to cluster together over time, along with the previous-round global model (S)
@@
The trajectories in FBPs enable the identification of different client behaviours during training (Figure 3). Figure 3 illustrates FBPs’ client trajectories over the last 15 rounds, where we introduced a different attack on each dataset (from left to right: Inverted-loss, Crafted-noise, and Inverted-gradient attacks). Each dataset was configured with five non-IID clients and one attacker (red). In the Synthetic dataset, to highlight the visualization of honest client clusters, the two largest clients were subdivided into three (Client 2,3,4) and two smaller clients (Client 6,7), respectively, forming distinct clusters. Specifically, on the EBP for the Synthetic dataset, the attacker (Client 9) significantly deviates from the trajectories of other clients, aiming to disrupt the server model (S). On the CBP, clusters with similar data distributions converge, indicating similarity in decision boundaries and client training data. In Breast Cancer and small-MNIST, FBPs are also able to give insights on the type of attack. On the Breast Cancer EBP, Client 6 exhibits a random trajectory around the server trajectory, which indicates a Crafted-noise attack. In small-MNIST, both the CBP and EBP reveal that at each round the trajectory of the attacker consistently moves in the opposite direction to that of the server, which indicates an Inverted-gradient attack. These insights might be useful for FL users as they explain the nature of clients’ behaviours, allowing the identification of as various types of attacks. Additional visualizations are available in the Appendix B.3.
@@
FBPs allow the identification of clusters of clients (see Figure 3). In the Synthetic dataset, CBPs’s clusters reflect client-specific data distributions (Client 1,2-3-4,5,6-7,8 and Attacker 9). Clients sharing similar data distributions tend to cluster closely in the CBP, which in the Synthetic dataset is structured with adjacent slices in the feature space (for example, Client 1 is positioned between Client 8 and Client 2). Similarly, in Breast Cancer and small-MNIST, two primary clusters are identifiable in the CBP: honest clients and attackers. This clustering capability is crucial for users aiming to comprehend client characteristics without direct access to data or models’ parameters, thereby informing strategic decisions in training a federated model. For instance, one might consider reducing the number of clients with identical data distributions to avoid redundancy and enhance training efficiency. Conversely, identifying distinct clusters in FBPs can be useful to maximise model performance (at the expense of the generalisation ability) by forming a Clustered FL system where independent federated models are trained using a subset of similar clients.